At the year end, the total assets of a business were £189,000 and total liabilities were £126,000. During the year, the owner withdrew £10,000 and introduced a further £18,000 in equity. The business began the year with total assets of £180,000 and total liabilities of £112,000. What is the profit(loss) for the year?
Answer #1

Beginning equity = Assets-Liabilities = 180000-112000 = 68000

Ending equity = 189000-126000 = 63000

Profit (loss) = 68000+18000-10000-63000 = -13000